# Preserving feasible space variational quantum algorithm for solving the uncapacitated facility location problem

### 18 Dec 2023 v2

@@ I Introduction
Variational Quantum Algorithms (VQAs) are a hot class of hybrid quantum-classical algorithms, with one of the core of ansatz design, which are expected to realize quantum advantages on Noisy Intermediate-Scale Quantum (NISQ) devices 17 ; 18 . Ansatz is generally divided into two types: problem-agnostic ansatz and problem-inspired ansatz. The structure of the problem-agnostic ansatz carries no information about the problem itself and is mostly suited for optimization problems, such as the Hardware-Efficient Ansatz (HEA) Hardware-efficient , which has the advantage of reducing the circuit depth as much as possible and being able to be implemented efficiently on a quantum chip. The specific structure of the problem-inspired ansatz typically depends on the task at hand, such as Quantum Approximation Optimization Algorithm (QAOA) QAOA and Quantum Alternating Operator Ansatz (QAOA+) 31 to solve the combinatorial optimization problems, which have also been applied to many problems maxcut2018 ; maxcut2021 ; mvcp2022 ; cc2022 ; 33 ; 34 ; 35 ; 36 ; 37 . HEA and QAOA need to search for the target solution in the Overall Hilbert Space, which may result in an invalid solution. Different from them, the mixed operators of QAOA+ preserves the feasible subspace, allowing the algorithm to search for target solutions in the Feasible Space (a subspace of the entire Hilbert space) of the constrained optimization problems 31 , resulting in the complete elimination of an invalid solution. Since QAOA+ has a significantly smaller search space, finding a solution is potentially more efficient.
@@
The UFLP Kuehn and Hamburger is a combinatorial optimization problem, which is one of the most important NP-hard problems in location theory Daskin2003 , with important applications in fields such as resource allocation Armas and Juan2018 , logistics and transportation Cura2010 , computer networks Lazic2010 , and computer vision Lazic2009 . However, there still exists a research gap in the field of quantum algorithms for addressing this particular issue. Intuitively, the QAOA+ seem to be a good choice for solving UFLP. Nevertheless, we find that the phase-separation operators of QAOA+ introduce lot of CNOT and $R_{Z}$ gates, thereby leading to increased circuit depth. Wang et al 34 have already pointed out that the phase-separation operators only change the phase, and numerical experiments demonstrated that discarding it hardly decreases the success probability. If the UFLP is solved using QAOA+, to facilitate the construction the mixed Hamiltonian, it needs to be transformed into a constrained optimization problem containing unconstrained variables (variables not present in the constraints), and we call it the Unconstrained-Variables Problem (UVP). For UVP, after discarding the phase-separation operator, only single-qubit rotating gates act on the unconstrained variables, lacking entanglement gates, resulting in the inability to traverse enough quantum states, which may lower the success probability of finding a solution. Therefore, QAOA+ is not suitable for such problems. How to design quantum algorithm with low circuit depth to solve UFLP efficiently is particularly important.
@@
As the mixed operators in QAOA+ preserve the feasible subspace and HEA reduces the circuit depth, we consider combining the advantages of both to design a quantum algorithm for UFLP. In this paper, we propose the Preserving Feasible Space-Variational Quantum Algorithm (PFS-VQA), which performs mixed operators and HEA act on different qubits respectively. Specifically, to address UFLP, it is transformed into a UVP in two steps. First, we transform the inequality constraint into equality constraint by introducing slack variables. Second, to facilitate the construction of a mixed Hamiltonian, a penalty function approach is adopted by including a constraint as part of the objective function, resulting in a UVP, where the feasible space is composed of bit strings with a fixed Hamming weight 1. To verify the efficiency of the algorithm, we simulate it for UFLP with 12 instances using MindQuantum 51 , showing that it has at least 54% higher success probability compared to QAOA+, QAOA, and HEA, and at least 75% lower circuit depth compared to QAOA+ and QAOA. Moreover, our algorithm reduces the number of CNOT gates and parameter gates by at least 59% and 53% compared to QAOA+ and QAOA, respectively.

### 20 Dec 2023 v3

@@ I Introduction
Variational Quantum Algorithms (VQAs) are a hot class of hybrid quantum-classical algorithms, promising to realize quantum advantages on Noisy Intermediate-Scale Quantum (NISQ) devices 17 ; 18 , and one of the core is the design of ansatz. Ansatz is generally divided into two types: problem-agnostic ansatz and problem-inspired ansatz. The structure of the problem-agnostic ansatz carries no information about the problem itself and is mostly suited for optimization problems, such as the Hardware-Efficient Ansatz (HEA) Hardware-efficient , which has the advantage of reducing the circuit depth as much as possible and being able to be implemented efficiently on a quantum chip. The specific structure of the problem-inspired ansatz typically depends on the task at hand, such as Quantum Approximation Optimization Algorithm (QAOA) QAOA and Quantum Alternating Operator Ansatz (QAOA+) 31 to solve the combinatorial optimization problems, which have been applied to many problems maxcut2018 ; maxcut2021 ; mvcp2022 ; cc2022 ; 33 ; 34 ; 35 ; 36 ; 37 . HEA and QAOA need to search for the target solution in the Overall Hilbert Space, which may result in an invalid solution. Different from them, the mixed operators of QAOA+ preserves the feasible subspace, allowing the algorithm to search for target solutions in the Feasible Space (a subspace of the entire Hilbert space) of the constrained optimization problems 31 , resulting in the complete elimination of an invalid solution. Since QAOA+ has a significantly smaller search space, finding a solution is potentially more efficient.
@@
The UFLP Kuehn and Hamburger is a combinatorial optimization problem, which is one of the most important NP-hard problems in location theory Daskin2003 , with important applications in fields such as resource allocation Armas and Juan2018 , logistics and transportation Cura2010 , and computer vision Lazic2009 . However, there still exists a research gap in the field of quantum algorithms for addressing this particular issue. Intuitively, the QAOA+ seem to be a good choice for solving UFLP. Nevertheless, we find that the phase-separation operators of QAOA+ introduce lot of CNOT and $R_{Z}$ gates, thereby leading to increased circuit depth. Wang et al 34 have already pointed out that the phase-separation operators only change the phase, and numerical experiments demonstrated that discarding it hardly decreases the success probability. If the UFLP is solved using QAOA+, to facilitate the construction the mixed Hamiltonian, it needs to be transformed into a constrained optimization problem containing unconstrained variables (variables not present in the constraints), and we call it the Unconstrained-Variables Problem (UVP). For UVP, after discarding the phase-separation operator, only single-qubit rotating gates act on the unconstrained variables, lacking entanglement gates, resulting in the inability to traverse enough quantum states, which may lower the success probability of finding a solution. Therefore, QAOA+ is not suitable for such problems. How to design quantum algorithm with low circuit depth to solve UFLP efficiently is particularly important.
@@
As the mixed operators in QAOA+ preserve the feasible subspace and HEA reduces the circuit depth, we consider combining the advantages of both to design a quantum algorithm for UFLP. In this paper, we propose the Preserving Feasible Space-Variational Quantum Algorithm (PFS-VQA), which performs mixed operators and HEA act on different qubits respectively. Specifically, to address UFLP, it is transformed into a UVP in two steps. First, we transform the inequality constraint into equality constraint by introducing slack variables. Second, to facilitate the construction of a mixed Hamiltonian, a penalty function approach is adopted by including a constraint as part of the objective function, resulting in a UVP, where the feasible space is composed of bit strings with a fixed Hamming weight 1. To verify the efficiency of the algorithm, we simulate it for UFLP with 12 instances using MindQuantum 51 , showing that it has at least 54% higher success probability and requires only about 125 iterations to converge to the optimum, with faster and better convergence compared to QAOA+, QAOA, and HEA. In addition, our algorithm is at least 75% lower circuit depth and reduces the number of CNOT gates and parameter gates by at least 59% and 53% compared to QAOA+ and QAOA, respectively.